# Towards a theory of dissipative Dark Matter I: the Born limit

### v1

@@ B.1 Contact diagram
$\sum_{k=1,4}\mathcal{M}_{c,k}$ $\simeq\left(\frac{g_{1}}{m_{1}}-\frac{g_{2}}{m_{2}}\right)\lambda_{S}\frac{\mu\left(\vec{v}_{i}-\vec{v}_{f}\right)\cdot\vec{e}}{\omega}+\mathcal{O}(v_{i}^{0}).$ (B.1)
Here, the polarization vectors are given in Coulomb gauge, $\epsilon^{\mu}(q)=(0,\mathbf{e})$ where $\mathbf{e}$ is one of the two transverse polarization three-vectors satisfying $\mathbf{q}\cdot\mathbf{e}=0$. The expression above is of order $\mathcal{O}(v_{i}^{-1})$, whereas dipole $\sim\mathcal{O}(v_{i}^{-3})$ and quadrupole $\sim\mathcal{O}(v_{i}^{-2})$ emission amplitudes are of lower order in velocity. We note that the associated squared amplitude, averaged (summed) over initial (final) spins, summed over polarizations, and angular averaged using (3.17), yields the energy differential cross section (3.26) and the energy loss rate (5.8), each multiplied by a factor $m_{V^{\prime}}^{4}/(4m_{1}m_{2})^{2}$ .
@@
For $S_{1}S_{2}\to S_{1}S_{2}\phi$,  $\tilde{S}_{1}\tilde{S}_{2}\to\tilde{S}_{1}\tilde{S}_{2}\phi$ similar conclusions on the relative size of quartic and trilinear couplings can be drawn with the replacement $g_{j}\to A_{j}/2m_{j}$ and $\mathbf{e}\to\mathbf{\bm{\hat{}}{\mathbf{q}}}$. The adopted condition (B.2) then again requires a hierarchy among the couplings of the dSIDM scenario that is hard to achieve in a perturbative regime, allowing us to neglect these contributions.
@@ B.2.1 Long-range mediated force
$\left\langle|\mathcal{M}_{\text{VIB}}|^{2}\right\rangle_{\mathbf{\bm{\hat{}}{\mathbf{q}}}}=A_{\phi}^{2}a^{2}_{1}a^{2}_{2}\frac{16m_{1}^{2}m_{2}^{2}}{\left(\mu^{2}|\mathbf{v}_{i}-\mathbf{v}_{f}|^{2}+m_{\phi}^{2}\right)^{4}},$ (B.3)
with $a_{j}=y_{j}$ for fermionic DM and $a_{j}=A_{j}/2m_{j}$ for scalar DM. In this form, similary to external leg emission, the corresponding differential cross-section for the emission of a soft scalar can be expressed as,
@@
Moreover, the stability of the scalar potential requires the inclusion of a quartic self-interaction term $\lambda_{\phi}\phi^{4}$ with $\lambda_{\phi}>0$, limiting the trilinear coupling as $A_{\phi}\lesssim m_{\phi}\sqrt{\lambda_{\phi}}$ [111]. In our NR power counting scheme, this translates to a coupling at most of order $A_{\phi}\sim\mathcal{O}(\mu v_{i}^{2})$. However, saturating this upper limit, combined with (B.5) violates the kinematic requirement $m_{\phi}<1/2\mu v_{i}^{2}$ for emission. We therefore require $A_{\phi}\ll\mu v_{i}^{2}$ which result, despite the velocity enhancement of (B.3), on processes that are tightly constrained by consistency arguments and, in conclusion, subleading to quadrupole emission.
@@
In the case of distinguishable particle scattering, the energy differential cross-section is obtained from (B.3), and is given by
$\omega\frac{d\sigma}{d\omega}=A_{\phi}^{2}\frac{a_{1}^{2}a_{2}^{2}\mu^{4}v_{i}^{2}}{48\pi^{3}}\sqrt{1-\frac{4m_{\phi}^{2}}{\mu^{2}v_{i}^{4}x^{2}}}\frac{(1-x)^{1/2}\left[3m_{\phi}^{4}+6m_{\phi}^{2}(x-2)\mu^{2}+\mu^{4}v_{i}^{4}(x-4)(3x-4)\right]}{\left(m_{\phi}^{4}+2m_{\phi}^{2}\mu^{2}v_{i}^{2}\left(x-2\right)+\mu^{4}v_{i}^{4}x^{2}\right)^{3}}.$ (B.6)
@@
with $\{a,s\}=\{A/2m_{S},0\}$ for scalar DM candidates and $\{y,1/2\}$ for fermionic DM candidates. The associated energy differential cross-sections are
@@ B.2.2 Short-range mediated force
A similar VIB process can also be studied in the heavy mediator limit. In that case, the interaction does not relies on the self-trilinear coupling $\frac{1}{3!}A_{\phi}\phi^{3}$, but rather on the mixed trilinear coupling between the light and heavy scalar mediators $\frac{1}{2}A^{\prime}_{\phi}\phi^{\prime 2}\phi$. In this setting, the coupling $A^{\prime}_{\phi}$ can be treated as a free parameter. This freedom allows one to study VIB for both massless and massive emission, since the emitted particle is now distinct from the heavy mediator.
@@
with $a^{\prime}_{j}=y^{\prime}_{j}$ for fermionic DM and $a^{\prime}_{j}=A^{\prime}_{j}/2m_{j}$. The corresponding energy differential cross-section reads
@@ C.1 Long-range mediation
$\langle|\mathcal{M}|^{2}\rangle_{\mathbf{\bm{\hat{}}{\mathbf{q}}}}$ $=\langle|\mathcal{M}_{\text{external leg}}|^{2}\rangle_{\mathbf{\bm{\hat{}}{\mathbf{q}}}}+\langle|\mathcal{M}_{\text{vertex}}|^{2}\rangle_{\mathbf{\bm{\hat{}}{\mathbf{q}}}}+\langle|\mathcal{M}_{\text{cross}}|^{2}\rangle_{\bm{\hat{}}{\mathbf{q}}}.$ (C.1)
The first term corresponds to the contribution from the external leg emission diagram, which dominates the dissipative process when $A_{1}/m_{1},\,A_{2}/m_{2}\gg\lambda_{1},\,\lambda_{2}$,
$\langle|\mathcal{M}_{\text{external leg}}|^{2}\rangle_{\mathbf{\bm{\hat{}}{\mathbf{q}}}}$ $=\frac{A_{{1}}^{2}}{m_{1}^{2}}\frac{A_{{2}}^{2}}{m_{2}^{2}}\left(\frac{A_{1}}{m_{1}^{2}}-\frac{A_{2}}{m_{2}^{2}}\right)\frac{\Delta m_{12}}{\mu^{2}|\mathbf{v}_{i}-\mathbf{v}_{f}|^{4}}\left[\frac{\Delta m_{12}}{4}\left(\frac{A_{1}}{m_{1}^{2}}-\frac{A_{2}}{m_{2}^{2}}\right)\right$. (C.2)
$\left.-\left(\frac{A_{1}}{m_{1}^{3}}+\frac{A_{2}}{m_{2}^{3}}\right)\frac{2m_{1}m_{2}}{3}\right]+A_{1}^{2}A_{2}^{2}\left(\frac{A_{1}}{m_{1}^{3}}+\frac{A_{2}}{m_{2}^{3}}\right)^{2}\frac{T_{1}+11T_{2}}{60\omega^{2}|\vec{v}_{i}-\vec{v}_{f}|^{4}},$
where $\Delta m_{12}=m_{1}-m_{2}$ and $T_{1,2}$ are defined in (4.14) and (4.15). The second term in (C.1) corresponds to the contribution from the vertex emission diagram, which becomes relevant when $A_{1}/m_{1},\,A_{2}/m_{2}\ll\lambda_{1},\,\lambda_{2}$,
$\langle|\mathcal{M}_{\text{vertex}}|^{2}\rangle_{\mathbf{\bm{\hat{}}{\mathbf{q}}}}=\left(\frac{A_{1}\lambda_{2}+A_{{2}}\lambda_{{1}}}{\mu^{2}|\vec{v}_{i}-\vec{v}_{f}|^{2}}\right)^{2}.$ (C.3)
Finally, $\langle|\mathcal{M}_{\text{cross}}|^{2}\rangle_{\bm{\hat{}}{\mathbf{q}}}$ contains the cross terms involving both types of couplings, which become subdominant in the limits $A_{1}/m_{1},\,A_{2}/m_{2}\gg\lambda_{1},\,\lambda_{2}$ or $A_{1}/m_{1},\,A_{2}/m_{2}\ll\lambda_{1},\,\lambda_{2}$,
$\langle|\mathcal{M_{\text{cross}}}|^{2}\rangle_{\bm{\hat{}}{\mathbf{q}}}=-\frac{A_{1}A_{2}\left(A_{1}\lambda_{2}+A_{2}\lambda_{1}\right)}{\mu^{3}|\mathbf{v}_{i}-\mathbf{v}_{f}|^{4}}\left[\frac{4}{3}\left(\frac{A_{1}}{m_{1}^{3}}+\frac{A_{2}}{m_{2}^{3}}\right)+\left(\frac{A_{1}}{m_{1}^{2}}-\frac{A_{2}}{m_{2}^{2}}\right)\frac{\Delta m_{12}}{m_{1}m_{2}}\right].$ (C.4)
@@
$\langle|\mathcal{M_{Q}}|^{2}\rangle_{\mathbf{\bm{\hat{}}{\mathbf{q}}}}=A_{1}^{2}A_{2}^{2}\left(\frac{A_{1}}{m_{1}^{3}}+\frac{A_{2}}{m_{2}^{3}}\right)^{2}\frac{\left[3\left(|\vec{v}_{f}|^{4}+|\vec{v}_{i}|^{4}\right)-|\vec{v}_{f}|^{2}|\vec{v}_{i}|^{2}\left((\hat{v}_{i}\cdot\hat{v}_{f})^{2}+5\right)\right]}{15\omega^{2}|\vec{v}_{i}-\vec{v}_{f}|^{4}}.$ (C.5)
@@
$\langle|\mathcal{E}_{{S\phi}}|^{2}\rangle_{\mathbf{\bm{\hat{}}{\mathbf{q}}}}$ $=\left(\frac{A_{1}}{m_{1}^{2}}-\frac{A_{2}}{m_{2}^{2}}\right)\frac{\Delta m_{12}\mu^{2}}{m_{1}m_{2}}\left[\frac{\Delta m_{12}}{4m_{1}m_{2}}\left(\frac{A_{1}}{m_{1}^{2}}-\frac{A_{2}}{m_{2}^{2}}\right)\right$. (C.6)
$\left.-\frac{2}{3}\left(\frac{A_{1}}{m_{1}^{3}}+\frac{A_{2}}{m_{2}^{3}}\right)\right]+\left(\frac{A_{{1}}}{m_{1}^{3}}+\frac{A_{{2}}}{m_{2}^{3}}\right)^{2}\frac{\mu^{4}\left(T_{1}+11T_{2}\right)}{60\omega^{2}}.$
@@ C.2 Short-range mediation
$\langle|\mathcal{M}_{\text{external leg}}|^{2}\rangle_{\mathbf{\bm{\hat{}}{\mathbf{q}}}}$ $=-\frac{C_{A_{j},\lambda_{S}}^{2}\mu^{2}\Delta m_{12}}{m_{\phi^{\prime}}^{4}m_{1}m_{2}}\left(\frac{A_{1}}{m_{1}^{2}}-\frac{A_{2}}{m_{2}^{2}}\right)\left[\frac{1}{6}\left(\frac{A_{1}}{m_{1}^{3}}+\frac{A_{2}}{m_{2}^{3}}\right)\right$. (C.7)
$\left.-\frac{\Delta m_{12}}{4m_{1}m_{2}}\left(\frac{A_{1}}{m_{1}^{2}}-\frac{A_{2}}{m_{2}^{2}}\right)\right]+C_{A_{j},\lambda_{S}}^{2}\left(\frac{A_{1}}{m_{1}^{3}}+\frac{A_{2}}{m_{2}^{3}}\right)^{2}\frac{\mu^{4}\left(T_{1}+2T_{2}\right)}{60\omega^{2}m_{\phi^{\prime}}^{4}},$
where $C_{A_{j},\lambda_{S}}=\left(A^{\prime}_{1}A^{\prime}_{2}-\lambda_{S}m_{\phi^{\prime}}^{2}\right)$. The second term of (C.1) corresponds to the contribution from the vertex emission diagram, and reads
$\langle|\mathcal{M}_{\text{vertex}}|^{2}\rangle_{\mathbf{\bm{\hat{}}{\mathbf{q}}}}=\left(\frac{A^{\prime}_{1}\lambda^{\prime}_{2}+A^{\prime}_{{2}}\lambda^{\prime}_{{1}}}{m_{\phi^{\prime}}^{2}}\right)^{2}.$ (C.8)
Finally, $\langle|\mathcal{M}_{\text{cross}}|^{2}\rangle_{\bm{\hat{}}{\mathbf{q}}}$ contains the cross terms involving all four couplings, and is defined by
$\langle|\mathcal{M_{\text{cross}}}|^{2}\rangle_{\bm{\hat{}}{\mathbf{q}}}=$ $-\frac{C_{A_{j},\lambda_{S}}\mu}{3m_{\phi^{\prime}}^{4}}\left(A^{\prime}_{1}\lambda^{\prime}_{2}+A^{\prime}_{2}\lambda^{\prime}_{1}\right)\left[\left(\frac{A_{1}}{m_{1}^{3}}+\frac{A_{2}}{m_{2}^{3}}\right)+\left(\frac{A_{1}}{m_{1}^{2}}-\frac{A_{2}}{m_{2}^{2}}\right)\frac{3\Delta m_{12}}{m_{1}m_{2}}\right].$ (C.9)
@@
$\langle|\mathcal{M}_{\mathcal{Q}}|^{2}\rangle_{\mathbf{\bm{\hat{}}{\mathbf{q}}}}$ $=C_{A_{j},\lambda_{S}}^{2}\left(\frac{A_{1}}{m_{1}^{3}}+\frac{A_{2}}{m_{2}^{3}}\right)^{2}\frac{\mu^{4}\left[-2|\vec{v}_{f}|^{2}|\vec{v}_{i}|^{2}\left(2(\hat{v}_{i}\cdot\hat{v}_{f})^{2}+1\right)+3\left(|\vec{v}_{f}|^{4}+|\vec{v}_{i}|^{4}\right)\right]}{60\omega^{2}m_{\phi^{\prime}}^{4}}.$ (C.10)
@@
$\langle|\mathcal{E}_{{S\phi}}|^{2}\rangle_{\mathbf{\bm{\hat{}}{\mathbf{q}}}}$ $=-\frac{\mu^{2}\Delta m_{12}}{m_{1}m_{2}}\left(\frac{A_{1}}{m_{1}^{2}}-\frac{A_{2}}{m_{2}^{2}}\right)\left[\frac{1}{6}\left(\frac{A_{1}}{m_{1}^{3}}+\frac{A_{2}}{m_{2}^{3}}\right)\right$$\left.\right$. (C.11)
$\left.-\frac{\Delta m_{12}}{4m_{1}m_{2}}\left(\frac{A_{1}}{m_{1}^{2}}-\frac{A_{2}}{m_{2}^{2}}\right)\right]+\left(\frac{A_{1}}{m_{1}^{3}}+\frac{A_{2}}{m_{2}^{3}}\right)^{2}\frac{\mu^{4}\left(T_{1}+2T_{2}\right)}{60\omega^{2}}.$

### v2

@@ B.1 Contact diagram
$\sum_{k=1,4}\mathcal{M}_{c,k}$ $\simeq\left(\frac{g_{1}}{m_{1}}-\frac{g_{2}}{m_{2}}\right)\lambda_{S}\frac{\mu\left(\vec{v}_{f}-\vec{v}_{i}\right)\cdot\vec{e}^{*}}{\omega}+\mathcal{O}(v_{i}^{0}).$ (B.1)
Here, the polarization vectors are given in Coulomb gauge, $\varepsilon^{\mu*}(q)=(0,\mathbf{e}^{*})$ where $\mathbf{e}^{*}$ is one of the two transverse polarization three-vectors satisfying $\mathbf{q}\cdot\mathbf{e}^{*}=0$. The expression above is of order $\mathcal{O}(v_{i}^{-1})$, whereas dipole $\sim\mathcal{O}(v_{i}^{-3})$ and quadrupole $\sim\mathcal{O}(v_{i}^{-2})$ emission amplitudes are of lower order in velocity. We note that the associated squared amplitude, averaged (summed) over initial (final) spins, summed over polarizations, and angular averaged using (3.24), yields the energy-differential cross section (3.32) and the energy loss rate (5.8), setting $g^{\prime}_{j}=0$.
@@
For the processes $S_{1}S_{2}\to S_{1}S_{2}\phi$,  $\tilde{S}_{1}\tilde{S}_{2}\to\tilde{S}_{1}\tilde{S}_{2}\phi$ similar conclusions on the relative size of quartic and trilinear couplings can be drawn with the replacements $g_{j}\to A_{j}/2m_{j}$, $\mathbf{e}^{*}\to\mathbf{\bm{\hat{}}{\mathbf{q}}}$, and adding an overall minus sign. The adopted condition (B.2) then again requires a hierarchy among the couplings of the dSIDM scenario that is hard to achieve in a perturbative regime, allowing us to neglect these contributions.
@@ B.2.1 Long-range mediated force
$\frac{1}{d_{1}d_{2}}\left\langle|\mathcal{M}_{\text{VIB}}|^{2}\right\rangle_{\mathbf{\bm{\hat{}}{\mathbf{q}}}}=A_{\phi}^{2}a^{2}_{1}a^{2}_{2}\frac{16m_{1}^{2}m_{2}^{2}}{\left(\mu^{2}|\mathbf{v}_{i}-\mathbf{v}_{f}|^{2}+m_{\phi}^{2}\right)^{4}},$ (B.3)
with $a_{j}=y_{j}$ for fermionic DM and $a_{j}=A_{j}/2m_{j}$ for scalar DM. In this form, similarly to external leg emission, the corresponding differential cross section for the emission of a soft scalar can be expressed as,
@@
Moreover, the stability of the scalar potential requires the inclusion of a quartic self-interaction term $\lambda_{\phi}\phi^{4}$ with $\lambda_{\phi}>0$, limiting the trilinear coupling as $A_{\phi}\lesssim m_{\phi}\sqrt{\lambda_{\phi}}$ [118]. In our NR power counting scheme, this translates to a coupling at most of order $A_{\phi}\sim\mathcal{O}(\mu v_{i}^{2})$. However, saturating this upper limit, combined with (B.5) violates the kinematic requirement $m_{\phi}<1/2\mu v_{i}^{2}$ for emission. We therefore require $A_{\phi}\ll\mu v_{i}^{2}$ which results, despite the velocity enhancement of (B.3), on processes that are tightly constrained by consistency arguments and, in conclusion, subleading to quadrupole emission.
@@
In the case of distinguishable particle scattering, the energy-differential cross section is obtained from (B.3), and is given by
$\omega\frac{d\sigma}{d\omega}=A_{\phi}^{2}\frac{a_{1}^{2}a_{2}^{2}\mu^{4}x^{2}v_{i}^{4}}{48\pi^{3}}\sqrt{1-\frac{4m_{\phi}^{2}}{\mu^{2}v_{i}^{4}x^{2}}}\frac{(1-x)^{1/2}\left[3m_{\phi}^{4}-6m_{\phi}^{2}v_{i}^{2}\mu^{2}(x-2)+\mu^{4}v_{i}^{4}(x-4)(3x-4)\right]}{\left(m_{\phi}^{4}-2m_{\phi}^{2}\mu^{2}v_{i}^{2}\left(x-2\right)+\mu^{4}v_{i}^{4}x^{2}\right)^{3}}.$ (B.6)
@@
with $\{a,s\}=\{A/2m_{S},0\}$ for scalar DM candidates and $\{y,1/2\}$ for fermionic DM candidates. The associated energy-differential cross sections are
@@ B.2.2 Short-range mediated force
A similar VIB process can also be studied in the heavy mediator limit. In that case, the interaction does not rely on the self-trilinear coupling $\frac{1}{3!}A_{\phi}\phi^{3}$, but rather on the mixed trilinear coupling between the light and heavy scalar mediators $\frac{1}{2}A^{\prime}_{\phi}\phi^{\prime 2}\phi$. In this setting, the coupling $A^{\prime}_{\phi}$ can be treated as a free parameter. This freedom allows one to study VIB for both massless and massive emission, since the emitted particle is now distinct from the heavy mediator.
@@
with $a^{\prime}_{j}=y^{\prime}_{j}$ for fermionic DM and $a^{\prime}_{j}=A^{\prime}_{j}/2m_{j}$. The corresponding energy-differential cross section reads
@@ C.1 Long-range mediation
$\left\langle|\mathcal{M}|^{2}\right\rangle_{\mathbf{\bm{\hat{}}{\mathbf{q}}}}$ $=\left\langle|\mathcal{M}_{\text{external leg}}|^{2}\right\rangle_{\mathbf{\bm{\hat{}}{\mathbf{q}}}}+\left\langle|\mathcal{M}_{\text{vertex}}|^{2}\right\rangle_{\mathbf{\bm{\hat{}}{\mathbf{q}}}}+\left\langle|\mathcal{M}_{\text{cross}}|^{2}\right\rangle_{\bm{\hat{}}{\mathbf{q}}}.$ (C.1)
The first term corresponds to the contribution from the external leg emission diagram, which dominates the dissipative process when $(A_{1}/m_{1})^{2},\,(A_{2}/m_{2})^{2}\gg\lambda_{1},\,\lambda_{2}$,
$\left\langle|\mathcal{M}_{\text{external leg}}|^{2}\right\rangle_{\mathbf{\bm{\hat{}}{\mathbf{q}}}}$ $=\frac{A_{{1}}^{2}}{m_{1}^{2}}\frac{A_{{2}}^{2}}{m_{2}^{2}}\left(\frac{A_{1}}{m_{1}^{2}}-\frac{A_{2}}{m_{2}^{2}}\right)\frac{\Delta m_{12}}{\mu^{2}|\mathbf{v}_{i}-\mathbf{v}_{f}|^{4}}\left[\frac{\Delta m_{12}}{4}\left(\frac{A_{1}}{m_{1}^{2}}-\frac{A_{2}}{m_{2}^{2}}\right)\right$. (C.2)
$\left.+\left(\frac{A_{1}}{m_{1}^{3}}+\frac{A_{2}}{m_{2}^{3}}\right)\frac{2m_{1}m_{2}}{3}\right]+A_{1}^{2}A_{2}^{2}\left(\frac{A_{1}}{m_{1}^{3}}+\frac{A_{2}}{m_{2}^{3}}\right)^{2}\frac{T_{1}+11T_{2}}{60\omega^{2}|\vec{v}_{i}-\vec{v}_{f}|^{4}},$
where $\Delta m_{12}=m_{1}-m_{2}$ and $T_{1,2}$ are defined in (4.14) and (4.15). The second term in (C.1) corresponds to the contribution from the vertex emission diagram, which becomes relevant when $(A_{1}/m_{1})^{2},\,(A_{2}/m_{2})^{2}\ll\lambda_{1},\,\lambda_{2}$,
$\left\langle|\mathcal{M}_{\text{vertex}}|^{2}\right\rangle_{\mathbf{\bm{\hat{}}{\mathbf{q}}}}=\left(\frac{A_{1}\lambda_{2}+A_{{2}}\lambda_{{1}}}{\mu^{2}|\vec{v}_{i}-\vec{v}_{f}|^{2}}\right)^{2}.$ (C.3)
Finally, $\langle|\mathcal{M}_{\text{cross}}|^{2}\rangle_{\bm{\hat{}}{\mathbf{q}}}$ contains the cross terms involving both types of couplings, which become subdominant in the limits $(A_{1}/m_{1})^{2},\,(A_{2}/m_{2})^{2}\gg\lambda_{1},\,\lambda_{2}$ or $(A_{1}/m_{1})^{2},\,(A_{2}/m_{2})^{2}\ll\lambda_{1},\,\lambda_{2}$,
$\left\langle|\mathcal{M_{\text{cross}}}|^{2}\right\rangle_{\bm{\hat{}}{\mathbf{q}}}=-\frac{A_{1}A_{2}\left(A_{1}\lambda_{2}+A_{2}\lambda_{1}\right)}{\mu^{3}|\mathbf{v}_{i}-\mathbf{v}_{f}|^{4}}\left[\frac{4}{3}\left(\frac{A_{1}}{m_{1}^{3}}+\frac{A_{2}}{m_{2}^{3}}\right)+\left(\frac{A_{1}}{m_{1}^{2}}-\frac{A_{2}}{m_{2}^{2}}\right)\frac{\Delta m_{12}}{m_{1}m_{2}}\right].$ (C.4)
@@
$\left\langle|\mathcal{M_{Q}}|^{2}\right\rangle_{\mathbf{\bm{\hat{}}{\mathbf{q}}}}=A_{1}^{2}A_{2}^{2}\left(\frac{A_{1}}{m_{1}^{3}}+\frac{A_{2}}{m_{2}^{3}}\right)^{2}\frac{\left[3\left(|\vec{v}_{f}|^{4}+|\vec{v}_{i}|^{4}\right)-|\vec{v}_{f}|^{2}|\vec{v}_{i}|^{2}\left((\hat{v}_{i}\cdot\hat{v}_{f})^{2}+5\right)\right]}{15\omega^{2}|\vec{v}_{i}-\vec{v}_{f}|^{4}}.$ (C.5)
@@
$\left\langle|\mathcal{E}_{{S\phi}}|^{2}\right\rangle_{\mathbf{\bm{\hat{}}{\mathbf{q}}}}$ $=\left(\frac{A_{1}}{m_{1}^{2}}-\frac{A_{2}}{m_{2}^{2}}\right)\frac{\Delta m_{12}\mu^{2}}{m_{1}m_{2}}\left[\frac{\Delta m_{12}}{4m_{1}m_{2}}\left(\frac{A_{1}}{m_{1}^{2}}-\frac{A_{2}}{m_{2}^{2}}\right)\right$. (C.6)
$\left.+\frac{2}{3}\left(\frac{A_{1}}{m_{1}^{3}}+\frac{A_{2}}{m_{2}^{3}}\right)\right]+\left(\frac{A_{{1}}}{m_{1}^{3}}+\frac{A_{{2}}}{m_{2}^{3}}\right)^{2}\frac{\mu^{4}\left(T_{1}+11T_{2}\right)}{60\omega^{2}}.$
@@ C.2 Short-range mediation
$\left\langle|\mathcal{M}_{\text{external leg}}|^{2}\right\rangle_{\mathbf{\bm{\hat{}}{\mathbf{q}}}}$ $=\frac{m_{1}m_{2}C_{A^{\prime}_{j},\lambda_{S}}^{2}\mu^{2}\Delta m_{12}}{m_{\phi^{\prime}}^{4}}\left(\frac{A_{1}}{m_{1}^{2}}-\frac{A_{2}}{m_{2}^{2}}\right)\left[\frac{1}{3}\left(\frac{A_{1}}{m_{1}^{3}}+\frac{A_{2}}{m_{2}^{3}}\right)\right$. (C.7)
$\left.+\frac{\Delta m_{12}}{4m_{1}m_{2}}\left(\frac{A_{1}}{m_{1}^{2}}-\frac{A_{2}}{m_{2}^{2}}\right)\right]+m_{1}^{2}m_{2}^{2}C_{A^{\prime}_{j},\lambda_{S}}^{2}\left(\frac{A_{1}}{m_{1}^{3}}+\frac{A_{2}}{m_{2}^{3}}\right)^{2}\frac{\mu^{4}\left(T_{1}+2T_{2}\right)}{60\omega^{2}m_{\phi^{\prime}}^{4}},$
where $C_{A^{\prime}_{j},\lambda_{S}}=(A^{\prime}_{1}/m_{1})(A^{\prime}_{2}/m_{2})-\lambda_{S}m_{\phi^{\prime}}^{2}/(m_{1}m_{2})$. The second term of (C.1) corresponds to the contribution from the vertex emission diagram, and reads
$\left\langle|\mathcal{M}_{\text{vertex}}|^{2}\right\rangle_{\mathbf{\bm{\hat{}}{\mathbf{q}}}}=\left(\frac{A^{\prime}_{1}\lambda^{\prime}_{2}+A^{\prime}_{{2}}\lambda^{\prime}_{{1}}}{m_{\phi^{\prime}}^{2}}\right)^{2}.$ (C.8)
Finally, $\left\langle|\mathcal{M}_{\text{cross}}|^{2}\right\rangle_{\bm{\hat{}}{\mathbf{q}}}$ contains the cross terms involving all four couplings, and is defined by
$\left\langle|\mathcal{M_{\text{cross}}}|^{2}\right\rangle_{\bm{\hat{}}{\mathbf{q}}}=$ $-\frac{2m_{1}m_{2}C_{A^{\prime}_{j},\lambda_{S}}\mu}{3m_{\phi^{\prime}}^{4}}\left(A^{\prime}_{1}\lambda^{\prime}_{2}+A^{\prime}_{2}\lambda^{\prime}_{1}\right)\left[\left(\frac{A_{1}}{m_{1}^{3}}+\frac{A_{2}}{m_{2}^{3}}\right)+\left(\frac{A_{1}}{m_{1}^{2}}-\frac{A_{2}}{m_{2}^{2}}\right)\frac{3\Delta m_{12}}{2m_{1}m_{2}}\right].$ (C.9)
@@
$\left\langle|\mathcal{M}_{\mathcal{Q}}|^{2}\right\rangle_{\mathbf{\bm{\hat{}}{\mathbf{q}}}}$ $=m_{1}^{2}m_{2}^{2}C_{A^{\prime}_{j},\lambda_{S}}^{2}\left(\frac{A_{1}}{m_{1}^{3}}+\frac{A_{2}}{m_{2}^{3}}\right)^{2}\frac{\mu^{4}\left[-2|\vec{v}_{f}|^{2}|\vec{v}_{i}|^{2}\left(2(\hat{v}_{i}\cdot\hat{v}_{f})^{2}+1\right)+3\left(|\vec{v}_{f}|^{4}+|\vec{v}_{i}|^{4}\right)\right]}{60\omega^{2}m_{\phi^{\prime}}^{4}}.$ (C.10)
@@
$\left\langle|\mathcal{E}_{{S\phi}}|^{2}\right\rangle_{\mathbf{\bm{\hat{}}{\mathbf{q}}}}$ $=\frac{\mu^{2}\Delta m_{12}}{m_{1}m_{2}}\left(\frac{A_{1}}{m_{1}^{2}}-\frac{A_{2}}{m_{2}^{2}}\right)\left[\frac{1}{3}\left(\frac{A_{1}}{m_{1}^{3}}+\frac{A_{2}}{m_{2}^{3}}\right)\right$$\left.\right$. (C.11)
$\left.+\frac{\Delta m_{12}}{4m_{1}m_{2}}\left(\frac{A_{1}}{m_{1}^{2}}-\frac{A_{2}}{m_{2}^{2}}\right)\right]+\left(\frac{A_{1}}{m_{1}^{3}}+\frac{A_{2}}{m_{2}^{3}}\right)^{2}\frac{\mu^{4}\left(T_{1}+2T_{2}\right)}{60\omega^{2}}.$